### v2

# Extrapolation of Machine-Learning Interatomic Potentials for Organic and Polymeric Systems

Abstract
Machine-Learning Interatomic Potentials (MLIPs) have surged in popularity due to their promise of expanding the spatiotemporal scales possible for simulating molecules with high fidelity. The accuracy of any MLIP is dependent on the data used for its training; thus, for large molecules, like polymers, where accurate training data is prohibitively difficult to obtain, it becomes necessary to pursue non-traditional methods to construct MLIPs, many of which are based on constructing MLIPs using smaller, analogous chemical systems. However, we have yet to understand the limits to which smaller molecules can be used as a proxy for extrapolating macromolecular energetics. Here, we provide a “control study” for such experiments, exploring the ability of MLIP approaches to extrapolate between n=1-8 n-polyalkanes at identical conditions. Through Principal Covariates Classification, we quantitatively demonstrate how convergence in chemical environments between training and testing datasets coincides with an MLIP’s transferability. Additionally, we show how careful attention to the construction of an MLIP’s neighbor list can promote greater transferability when considering various levels of the energetic hierarchy. Our results establish a roadmap for how one can create transferable MLIPs for macromolecular systems without the prohibitive cost of constructing system-specific training data.
I Introduction
Figure 1: How well do ML potentials built on short-chain molecules generalize to their longer counterparts?
In molecular dynamics, the definition of “potential energy” is one of the most critical simulation components, determining both the relevance and performance of the resulting computational experiment. Thus, there are countless methods for determining the potential energy of a chemical system; these methods often play the balance between assumed accuracy and efficiency, with the more “accurate” methods focused on calculating or approximating the underlying electronic structure of a system, whereas the more “efficient” methods focus on simple mathematical functions that can be computed easily for many atoms across long length and time scales.
For polymeric systems, which typically contain molecules consisting of hundreds or thousands of atoms, the latter, more “efficient” approach, classical force fields, has been the necessary choice. In a classical force field, we assume that a potential energy function can be broken down into conceptually discrete components corresponding to bonds, torsions, dihedrals, and van der Waals interactions. This results in a simple, fixed-form potential energy function parameterized to fit a specific system, for which there are countless options[1], approaches, and strategies, to varying degrees of experimental fidelity and success. While these force-fields aim to be accurate, many struggle to generalize beyond their original reference system. Furthermore, they often require the use of specific “scale” or “fudge” factors to improve experimental agreement[2, 3]. As such, it is often hard to trust classical force fields “as-is,” and thus the results obtained from them are not considered representative until experimentally verified.
Recent years has seen an explosion of “machine-learned interatomic potentials” (MLIPs), which recast the “accurate” into “efficient” procedures, providing surrogate models to electronic structure calculations that can be computed on-the-fly[4, 5, 6, 7, 8, 9, 10, 11]. Such MLIPs not only provide accurate and fast energetic predictions, but several strategies have been utilized that improve the generalizability of these models[12, 13, 14]. Similar to traditional force-field fitting, MLIPs work by determining potential energy functions for molecular configurations from a set of training configurations. The power in these approaches is their flexibility in the potential energy functional form, alleviating the source of many errors in classical force fields and enabling the study of ionic transport[15, 16], condensed-phase phenomena[17, 18], and electronic properties[19].
A challenge in applying these approaches to macromolecular systems stems from the fact that MLIPs, particularly those built for a specific system, require high-quality, ab initio training data on which to build a machine-learning model. This, in itself, is prohibitive for macromolecular systems, where such calculations are less feasible. Thus, despite the many problems in polymer science that require a quantum chemical level of accuracy, specially-trained MLIPs are out of reach[20].
So, how do we harness these powerful methods for polymeric simulation, where it is difficult, if not infeasible, to construct the requisite underlying training data? “Universal” or “foundational” MLIPs (UMLIPs) have recently emerged as a solution, as they promise to yield accurate potential energies for a wide variety of systems from pre-trained models[21, 22, 23, 24]. These potentials, owing to their recency, have yet to be tested for polymeric systems. Yet, while UMLIPs are generally applicable to many systems, there will always exist cases where custom training will be required.
To create a custom MLIP for polymeric systems, a popular approach is molecular fragmentation[25], in which one performs ab-initio calculations on small molecular fragments and combines them to calculate the energy of larger molecules. The molecular fragmentation approaches have been utilized within MLIP frameworks such as QRNN[26, 27] and GEMS[28]. These combinations yielded a smooth PES that resulted in stable and accurate molecular dynamics simulations. However, while powerful, this approach still requires explicit classical simulation of the overarching system to generate fragments, for which FFs are prone to their standard pitfalls.
As many macromolecular molecules include consistent, repeated chemistries, we hypothesize that we can create an MLIP trained on smaller analogs of a macromolecule and extrapolate the energetics to longer systems, as posited in Fig. 1. To test this hypothesis, we curate a dataset specifically for this purpose – n=18n=1-8 linear alkanes simulated at similar conditions (i.e., phase, density, pressure, and temperature). By doing so, we can provide critical guidance to MLIP construction for polymeric systems, both using traditional approaches and molecular fragmentation.
Our analysis shows that extrapolation is possible once all relevant carbon and hydrogen environment statistics have converged within the training set. This result, while statistically demonstrated, is nuanced, as extrapolating pre-built potentials requires additional procedures beyond those typically practiced in the community. And thus, we end our discussion with practical steps how to extrapolate polymeric MLIPs.
II Methods
II.1 Data Generation
Each MLIP considered was trained on a single n=18n=1-8 $n$-polyalkane dataset. To limit confounding factors in our analyses, datasets were constructed at similar phase, pressure, and temperature conditions. Each dataset consisted of $N=64$ molecules, simulated at T=300T=300K with the volume corresponding to NIST fluid properties’ density for each alkane at 5 MPa and 300K[29]. At these temperature and pressure conditions, each alkane is in a liquid state, aside from methane, which is a supercritical liquid. Initial configurations of n=38n=3-8 polyalkanes were obtained from the Cambridge Structural Database[30]. For methane and ethane, initial configurations were generated using RDKit[31] on a cubic lattice with an appropriate volume according to NIST fluid properties’ data. Each initial configuration was relaxed to an unordered, liquid state using isobaric isothermal (NPT) molecular dynamics simulations at low pressure ($P=0.5$ MPa, T=300T=300K ) using density-functional tight binding (DFTB+)[32] and the mio Slater-Koster files[33]. These NPT simulations were performed until the appropriate density corresponding to NIST fluid properties’ data was obtained for each polyalkane. For $n=7,8$ datasets, further relaxation at large box volume (with $P=0.01$ MPa, T=300T=300K) was necessary in order to remove initialization artifacts. After these systems were randomized, LAMMPS[34] was used to repressurize these two datasets to 300K and 5 MPa with the OPLS-AA forcefield until they reached the appropriate volume [35]. With $n=8$, some initialization artifacts still remained, so we computed structure optimization in ASE using a BFGS algorithm and fmax=0.05 with a DFTB+ calculator using a Conjugated Gradient driver to anneal the system.
NVT simulations were then conducted at 300K. These production simulations were conducted using DFTB+, initially using a grid of (1, 1, 1) k-points, and were run until potential energies had converged. For our purposes, an equilibrium cutoff was defined as where the running average of the total energy did not fluctuate by more than 0.1 eV for 2,000 consecutive frames in the NVT simulation. We then selected 1000 frames from the subsequent simulation at random to form testing sets, and used farthest-point sampling[36, 37, 38] on Smooth Overlap of Atomic Positions (SOAP[4]) vectors to select a mutually-exclusive 10,000 frames to constitute our training sets. We recomputed the energies and forces for the training and testing datasets using higher grid-point interpolation (3, 3, 3) using a high-throughput workflow with signac[39] and signac-flow[40].
To construct values for inter- and intra-molecular energies and forces, we isolated the individual molecules into separate ASE-atoms objects and recomputed energies and forces without consideration of periodic boundary conditions. This yielded
$E_{intermolecular,\ bulk}=E_{total,\ bulk}-E_{intramolecular}$
where $E_{intramolecular}\equiv\sum_{i=1}^{64}E_{i}$, with $E_{i}$ the total energy of the $i^{th}$ single molecule.
II.2 MLIP Training
For every MLIP, we solely used the aforementioned “training” sets during construction, and trained each model as one normally would for a non-extrapolative task. When training solely on intramolecular quantities (Fig. S3), we down-selected from the 640,000 single-molecule frames to 10,000, choosing one random molecule from each training frame. Testing sets were kept consistent, with 64,000 single molecule frames in total.
II.2.1 MACE
The MACE[5] model (Higher‑Order Equivariant Message Passing Neural Network) combines equivariant graph neural network design with explicit many‑body message passing. Unlike most standard MPNNs that rely on only two‑body (pairwise) interactions and require many message‑passing layers to build complexity, MACE constructs higher‑body (up to four‑body) equivariant messages in a hierarchical fashion. This allows the network to achieve high representational power with just two message‑passing iterations, substantially reducing depth, improving parallelism, and maintaining scalability while reaching or surpassing state‑of‑the‑art accuracy on benchmarks such as rMD17, 3BPA, and AcAc[5]. By building hierarchical tensor products of lower‑order features $A^{(t)}_{i}$ and symmetrizing via generalized Clebsch–Gordan coupling, MACE efficiently implements these many‑body messages without exponential cost. This induces equivariant feature tensors $B^{(t)}_{i,\eta\nu,kLM}$, which are linearly combined (with weights dependent on element type and irreducible representation indices $k,L,M$) to produce message updates $m^{(t)}_{i,kLM}$. For our MACE-based potentials, we use the software v.0.3.13 and a standard set of model hyperparameters, as shown in the example script in App. A.
(a) Energy-fitting results of MACE MLIP for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
(b) Force-fitting results of MACE MLIPs for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
(c) The relationship between the shift in composition $(x_{C}^{\text{train}}-x_{C}^{\text{test}})$ and the shift in energies for extrapolated MLIPs. The models trained on the methane, ethane, and propane datasets show a divergence from this proportionality, as they are not necessarily considered “well-conditioned.”
(d) Shifted energy errors for MACE energies, accounting for the learnable shift in Fig. 2(c).
Figure 2: Results of MACE potential building on the total potential energy on n=18n=1-8 alkanes.
II.2.2 SOAP Analysis
Smooth Overlap of Atomic Positions (SOAP) vectors were computed using the featomic[41] and librascal [42] packages with an interaction cutoff of $5\AA$ (for Fig. 3, Fig. S4) and $7\AA$ (for Fig. 4, Fig. S5); the GTO basis set was computed with 8 radial components and 4 angular components. Principal Covariates Classification (PCovC, [43]) was computed using v0.3.2 of scikit-matter[36], using as $X$ the aforementioned SOAP vectors of the training set, a mixing parameter of $0.5$, a logistic regression classifier, and oligomer length $n$ as our class value. Each PCovC model was computed solely for like-environments, such as H-, CH2, or CH3 environments (Fig. 3, Fig. S4). Prior to fitting and analysis, feature vectors and energies were appropriately scaled and centered using scikit-matter[36] StandardFlexibleScaler. For regression (Fig. 4), we used scikit-learn[44] RidgeCV with cv=5 and a negative mean squared error scorer. 20 regularization constants ($\alpha$) were considered for RidgeCV using a logarithmic grid.
III Results
On first glance, the energy-fitting results of Fig. 2(a) tell a discouraging story – each training set is unable to predict the energies of the other polyalkanes, with a MAE well-above chemical accuracy (which we define as 1meV/atom[45]). However, the force results (Fig. 2(b)) imply something much different, in line with our chemical intuition – at larger chain lengths, we can reasonably extrapolate the energetics of one system to the next (and even those below).
So, why do these two results diverge? A first intuition would be data centering – machine-learning potentials are typically trained to predict the fluctuations in molecular energetics, not their absolute value. In other words, the majority of MLIP methods and software first subtracts the mean of the training targets before determining weights and biases. This means that the predicted energy $E_{pred}$ is ostensibly
$E_{pred}=E_{MLIP}+b,$
where $b$ is a “baselining constant,” or the energetic values shared by all training points. This is verified by the strong colinearity between our predicted and true values, as shown in Fig. S2. A simple explanation of this “mean-shift” would be the shift in atomic energies,
$E_{atom}=n_{C}E_{C}+n_{H}E_{H}=N(x_{C}E_{C}+x_{H}E_{H}),$
which, when learning per-atom energies $E/N$, would yield
$b\equiv(x_{C}^{\text{train}}-x_{C}^{\text{test}})(E_{C}-E_{H}).$
However, this is not the case here, where the $E_{C}-E_{H}$ is on the order of 30eV (see App. A), leading to a much larger value than the observed shift. This presents an issue for extrapolating energetics from one system to another, where the only information known a priori may be the difference in underlying chemistry, and where “transfer learning” may be infeasible due to the lack of underlying training data. Thus, we first determine if $b$ is a linear function of our composition ($x_{C}$, $x_{H}$), as this would still allow for the extrapolation of one system to another without prior computation of energetics or configurations.
For models trained on n=48n=4-8 alkanes, there exists a strong linear proportionality (Pearson correlation coefficient = 0.9997, Fig. 2(c)) between the compositional shift and the mean-shift of the datasets. This suggests that these mean-shifts are learnable parameters, even if not obviously grounded in physical values such as $E_{atom}$. To test this, we determined a slope to the line in Fig. 2(c) via non-regularized linear regression, omitting from training the values obtained from models built on methane (blue), ethane (light blue), and propane (green), which are considered ill-conditioned models. Accounting for the large co-linearity of the mean-shift and compositional change, many of the error predictions are improved, yet there exist some residual errors (Fig. 2(d)). This suggests that the change in conformational distribution contributes non-negligibly to the mismatch in energies.
However, the results of Fig. 2(b) still demonstrate that there exist two critical lengths of polyalkane for extrapolation: 1) butane, where we see a general reduction in error magnitude from 20-30meV/Å to 3-6meV/Å (omitting methane), and 2) hexane, after which we see little improvement in force error with additional molecular units. The change in errors from propane to butane follows our chemical intuition – prior to this chain length, there is no sampling of the dihedral rotation possible in these molecules, and thus any system (n=48n=4-8) that contains these interactions would be ill-represented by such models.
But what makes hexane special? To understand this, we employ hybrid supervised-unsupervised analyses on the underlying environments within the training sets, focusing on the three-body interactions of the CH2 terms, which demonstrate the greatest statistical difference from dataset to dataset (see Fig. S4 for further analyses). We do so by employing a preceding technology to MACE descriptors, the Smooth Overlap of Atomic Positions (SOAP[4]), which are deterministic in their description of atomic environments (as opposed to MACE descriptors, whose higher-order correlations are determined through the message-passing component of the MACE training). We consider only the descriptors based on the first layer of the MACE perceptive field (5.0Å) that will be most heavily weighted in the constructed regression. We pass the SOAP vectors through Principal Covariates Classification (PCovC)[43, 46, 36], which computes the two-dimensional projection that both retains PCA-like variance and best separates the data based on assigned labels. In this context, PCovC can be applied to identify the environments that distinguish one training set from another (by using the training set identities as our classification labels). As is shown in Fig.  S4, a PCA projection is unable to yield such delineation in similar datasets.
Figure 3: Distribution of CH2 environments for the different training sets, mapped Principal Covariates Classification (PCovC[43]). The map demonstrates the convergence of environments with increasing $n$.
From this analysis, we can see that the convergence in MLIP error corresponds with the sampling of “E” environments with hexane, and the addition of “F” environments within heptane and octane leads to diminishing returns in terms of force approximation. This demonstrates that determining the minimal chain length required for MLIP construction (or molecular fragmentation[25, 26, 27, 28]) amounts to determining the convergence of the majority of molecular environments, a task that can be accomplished via classification models or similar unsupervised tasks on single-molecule data.
However, predicting the total energies and forces may not be sufficient for a usable potential model, as the hierarchical nature of interactions within these systems decides much of their fundamental behavior[13]. Here, we consider potential energies emerging from three distinct classes of interactions:
$E=E_{atom}+E_{intramolecular}+E_{intermolecular},$
as these terms have significantly different magnitudes, difficulties in calculating, and difficulties in terms of “learnability.” $E_{atom}$ is the largest component but arguably trivial to compute. We can group $E_{atom}$ with intramolecular interactions to obtain an $\tilde{E}_{intramolecular}=E_{atom}+E_{intramolecular}$ by sampling single molecule configurations; this is likewise a large component of $E$. As shown in Fig. S3, extrapolating $\tilde{E}_{intramolecular}$ follows the same lessons learned in predicting $E$ – both in terms of the necessary molecule lengths for extrapolation and the ML baselining leading to divergent narratives for energy- and force-fitting.
The element of this energetic expansion that requires the greatest nuance is the intermolecular interactions $E_{intermolecular}$, which are comparatively weak, longer-range, and computationally expensive to calculate using methods from quantum chemistry[47]. And, when training on total energies, the resultant errors often overshadow these small contributions[13], losing interactions that are decisive in many critical scientific problems.
MACE itself (v.0.3.13) is not natively suited for directly learning these interactions, nor, would we argue, are atom-centered expansions that implicitly or explicitly focus on the nearby atoms in constructing their MLIP. To demonstrate this, we again turn to the precursor for MACE[5], SOAP[4], and show how careful attention to weighting interactions can yield better fidelity for intermolecular models.
Following the same procedures in terms of training/testing splits, we construct cross-validated Ridge Regressions on SOAP representations for the same n=18n=1-8 $n$-polyalkane datasets used for MACE MLIP construction. SOAP vectors were calculated with a cutoff of 7Å for atomic environments and averaged for each frame such that for the 10,000 training frames and 1,000 testing frames, we have 10,000 and 1,000 corresponding SOAP vectors; we refer to these initial SOAP calculations as the “total” SOAP vectors, or $X^{\text{total}}$.
(a) Regressions using total SOAP vectors $X^{\text{total}}$.
(b) Regressions using the far-sighted SOAP vectors $X^{fs}$.
Figure 4: Results of SOAP-Ridge MLIPs trained on intermolecular potential energy using different SOAP representations .
Similar to the initial results from MACE in Fig. 2(a), the results of using a $E_{\text{intermolecular}}\approx X^{\text{total}}w$, shown in Fig. 4(a), are discouraging. Extrapolation, with a threshold MAE of 1 meV/atom [45], is only achieved three times (by hexane on heptane, and heptane on hexane/octane). However, we can adopt lessons from learning the lattice energy of molecular crystals[48] – SOAP representations, and representations that are implicitly or explicitly “near-sighted”[5, 49, 50] will, by design, consider the contributions of nearby atoms more strongly than those further away. This will overshadow the intermolecular environments which contribute to $E_{intermolecular}$, limiting our accuracy in extrapolation.
We can thus construct a “far-sighted” SOAP vector
$X^{fs}=X^{total}-\frac{1}{64}\sum_{i=1}^{64}X_{i}$
where $X_{i}$ is the averaged SOAP vector of the atomic environments in the $i$th molecule within the bulk frame. Effectively, this representation re-weights the feature space by reducing (or eliminating) the influence of contributions from intramolecular environments on the SOAP vector. As shown in Fig. 4(b), using this far-sighted representation results in a significant improvement in the number of extrapolative regimes achieved by SOAP-Ridge MLIPs in comparison to the results in Fig. 4(a). Similar trends are observed as in Fig. 2(b): steep improvements in MAEs are observed between ethane and propane, and once again from pentane to hexane. This follows the convergence in chemical environments (Fig. 3, Fig. S4). The results shown in Fig. 4 demonstrate the importance of carefully considering neighbor list construction when making MLIPs for different levels of the energetic hierarchy.
IV Conclusions
In this study, we have demonstrated key behaviors of MLIPs that can be applied to build predictive models for oligomeric systems, with implications on MLIP extrapolation for macromolecular and polymeric materials. First, we demonstrate that, in systems with similar chemistries, extrapolation of forces and energies is possible, provided that those environments present in the target molecule are properly sampled within the smaller analogs. Second, moving from one chemistry to another results in a learnable shift in energies; this can be largely ameliorated by predicting the shift in probability distributions with shifting composition. Finally, we demonstrated a powerful lesson about the hierarchy of energies – intermolecular energetics can be largely extrapolated by shifting the attention in model construction away from nearby neighbors. Given the relative difficulty of computing intra- versus intermolecular energetics, this suggests that the most difficult calculations, those relating the interactions of one molecule to another, are the easiest to translate from one system to another of similar chemistry. These results establish a blueprint for polymeric materials, suggesting that many problems are within reach with careful construction of potential energy models.
V Research Funding
This project was funded by the Wisconsin Alumni Research Fund (RKC) and by NSF through the University of Wisconsin Materials Research Science and Engineering Center (DMR-2309000, AL). NH acknowledges support from the Michael and Virginia Conway Professorship. The funders played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
VI Data Availability
The datasets generated and/or analysed during the current study are available in the Zenodo repository at Hooven et al. [51] [DOI: 10.5281/zenodo.17227874].
VII Code Availability
The underlying code used in this study are available via open-source software through GitHub, namely MACE (https://github.com/ACEsuit/mace[5]), scikit-matter (https://github.com/scikit-learn-contrib/scikit-matter[36]), featomic (https://github.com/metatensor/featomic[41]), and librascal (https://github.com/lab-cosmo/librascal.git[42]).
VIII Author Contributions
Project conceptualization (NH, RC), data curation (NH, AL), formal analysis (NH, RC), methodology (NH, RC), supervision (AL, RC), writing (NH, AL, RC).
IX Competing Interests
All authors declare no financial or non-financial competing interests.
Appendix A Additional methodological details
MACE MLIPs were constructed using the following command and MACE (v.0.3.13):
Listing 1: Training script for MACE MLIPs
raw_dir=\$(python -c "import mace; print(mace.__file__)"); dir=\${raw_dir:0:-12}
python \$dir/cli/run_train.py \
--name="MACE_model" \
--train_file="\$1" \
--valid_fraction=0.05 \
--config_type_weights=’{"Default":1.0}’ \
--E0s=’{1:-6.492647589968434, 6:-38.054950840332474}’ \
--model="MACE" \
--hidden_irreps=’64x0e + 64x1o’ \
--r_max=5.0 \
--batch_size=1 \
--max_num_epochs=1500 \
--energy_key="energy_eV" \
--forces_key="forces" \
--charges_key="charges" \
--amsgrad \
--restart_latest \
--device=cuda \
--patience=20 \
--default_dtype="float32" \
SOAP vectors for SOAP-Ridge MLIPs were computed using the following hyperparameters with librascal (v.0.0.1):
Listing 2: Hyperparameters of SOAP vectors for SOAP-Ridge MLIPs
"interaction_cutoff": 7,
"max_radial": 8,
"max_angular": 4,
"gaussian_sigma_constant": 0.3,
"gaussian_sigma_type": "Constant",
"cutoff_smooth_width": 0.5,
"radial_basis": "GTO",
"cutoff_function_type": "RadialScaling",
"cutoff_function_parameters": {
"rate": 1.5,
"exponent": 3.0,
"scale": 2.0
Appendix B Supporting figures to the text
(a) Alternative representation of Fig. 2(a).
(b) Alternative representation of Fig. 2(b).
Figure S1: Additional visualizations of MACE models to predict total energy and forces.
Figure S2: Map of the coefficient of determination between true and predicted values (represented as the logarithm of $1-R^{2}$)
(a) Energy-fitting results of MACE MLIP for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
(b) Force-fitting results of MACE MLIPs for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
(c) The relationship between the shift in composition $(x_{C}^{\text{train}}-x_{C}^{\text{test}})$ and the shift in energies for extrapolated MLIPs. The models trained on the methane, ethane, and propane datasets show a divergence from this proportionality, as they are not necessarily considered “well-conditioned.” Marker denotes the testing set, and color denotes the training set.
(d) Shifted energy errors for MACE energies, accounting for the learnable shift in Fig. 2(c).
Figure S3: Analogous to Fig. 2, results of predicting the intramolecular potential energy ($\tilde{E}_{intramolecular}=E_{intramolecular}+E_{atom}$) on n=18n=1-8 alkanes.
(a) Maps of H environments. At a perceptive length of 5Å, we see a convergence of hydrogen environments once reaching pentane.
(b) Maps of CH2 environments. At a perceptive length of 5Å, we see a convergence of CH2 environments once reaching hexane.
(c) Maps of CH3 environments. At a perceptive length of 5Å, we see a convergence of CH3 environments once reaching butane.
Figure S4: Analogous to Fig. 3, maps of different environments for the different training sets, via Principal Components Analysis (left) and Principal Covariates Classification (PCovC[43], right)
(a) Relative errors of SOAP-Ridge potentials trained on $X^{total}$ and intermolecular energies.
(b) Relative errors of SOAP-Ridge potentials trained on $X^{fs}$ and intermolecular energies.
Figure S5: Analogous to Fig. 4, heatmaps of relative errors for different training and testing sets calculated using the mean-absolute-error in meV/atom for testing sets divided by the interpolative error (error of the training alkane on its testing set).