### v3

# Extrapolation of Machine-Learning Interatomic Potentials for Organic and Polymeric Systems

Abstract
Machine-Learning Interatomic Potentials (MLIPs) have surged in popularity due to their promise of expanding the spatiotemporal scales possible for simulating molecules with high fidelity. The accuracy of any MLIP is dependent on the data used for its training; thus, for large molecules like polymers and biomolecules, where ab initio training data is prohibitively difficult to obtain, it becomes necessary to use smaller, analogous chemical systems to construct MLIPs. Here, we perform a control study using n=18n=1-8 linear alkanes to determine when MLIPs trained on small molecules can accurately extrapolate to longer chains and more complex architectures. By combining MLIP performance analysis with environment‑resolved SOAP descriptors and Principal Covariates Classification, we show that reliable extrapolation emerges precisely when the relevant local chemical environments have converged between training and target systems. We further demonstrate that careful construction of the neighbor list substantially improves the learnability of intermolecular energetics, the component most critical for polymeric behavior. Together, these results provide a practical, data‑driven blueprint for designing transferable MLIPs for macromolecular materials, whether built from bespoke training data or adapted from emerging universal MLIP frameworks.
Introduction
In molecular simulation, the choice of potential energy model determines both the accuracy of predicted behavior and the computational time scales that can be achieved. Classical force fields have played a foundational role in polymer science for decades, enabling the study of thermodynamics, structure, and dynamics across a wide variety of molecular and polymeric systems. Their conceptual clarity, expressing molecular energetics through well‑defined bonded and nonbonded terms, combined with their computational efficiency, continues to make them indispensable for simulations that require long time scales, large system sizes, or broad chemical screening.
At the same time, advances in electronic‑structure theory, machine learning, and high‑performance computing have created new opportunities at the interface of accuracy and efficiency. Machine‑learned interatomic potentials (MLIPs) offer one such opportunity: by learning from quantum‑chemical reference data, they can capture complex many‑body interactions without requiring an explicit decomposition into fixed functional forms[2, 4, 6, 11, 5, 38, 36, 37]. MLIPs have recently demonstrated strong performance across problems ranging from ionic transport[49, 19], condensed-phase phenomena[10, 52], and electronic properties[23], motivating interest in when and how these models can complement existing force‑field approaches–particularly for systems where non-additive electronic effects or chemical heterogeneity play an important role.
However, using MLIPs for large molecules such as polymers or biomolecules remains challenging. High‑quality quantum chemical data are difficult to obtain for such macromolecules[50], and transferability across chain length or architecture is not guaranteed or well-understood. This has led to increasing interest in strategies where smaller, chemically related molecules act as surrogates for constructing potentials that are then applied to larger targets[29, 41, 48]. Although intuitive, the limits of such extrapolation have not been systematically quantified.
Universal or “foundational” machine-learned interatomic potentials (UMLIPs) have recently emerged as a promising direction, offering broad chemical coverage from large, diverse training sets[40, 3, 45, 51]. These models reduce the need for system‑specific quantum data and have shown strong performance across many small‑molecule and materials benchmarks. However, due to their recency, their behavior for macromolecular systems remains largely untested. Moreover, even when UMLIPs provide a reliable baseline, fine‑tuning or system‑specific adaptation is often needed for application-specific studies. And thus, understanding which molecular environments must be represented is still quintessential, regardless of whether one uses a bespoke MLIP or a UMLIP.
In this work, we provide a control study of this question. Using linear alkanes with n=18n=1-8 as a model system, we examine when MLIPs trained exclusively on small molecules can accurately extrapolate to longer chains and more complex architectures. By combining MLIP performance analysis with environment‑level descriptors and classification tools, we identify the molecular environments that govern extrapolation accuracy and show how neighbor‑list design influences the learnability of inter‑ versus intramolecular interactions. These results form a practical blueprint for building transferable MLIPs for macromolecular systems, in a way that complements, and in some cases extends, the capabilities of traditional force‑field approaches.
Methods
Data generation
Training sets
Each MLIP was trained on a single n=18n=1-8 $n$-polyalkane dataset. To limit confounding factors in our analyses, datasets were constructed at similar phase, pressure, and temperature conditions. Each dataset consisted of $N=64$ molecules, simulated at T=300T=300K with the volume corresponding to NIST fluid properties’ density for each alkane at 5 MPa and 300K[39]. At these temperature and pressure conditions, each alkane is in a liquid state, aside from methane, which is a supercritical liquid. Initial configurations of n=38n=3-8 polyalkanes were obtained from the Cambridge Structural Database[24]. For methane and ethane, initial configurations were generated using RDKit[35] on a cubic lattice with the appropriate volume. Each initial configuration was relaxed to an unordered, liquid state using isobaric-isothermal (NPT) molecular dynamics simulations at low pressure ($P=0.5$ MPa, T=300T=300K). These NPT simulations were performed with a Berendsen barostat and thermostat, a timestep of 0.2 femtoseconds as well as periodic boundary conditions, and were run until the reference experimental density was obtained for each polyalkane.
For $n=7,8$ datasets, further relaxation at large box volume (with $P=0.01$ MPa, T=300T=300K) and geometry optimization was necessary in order to remove initialization artifacts. After these systems were randomized, we repressurized these two datasets to 300K and 5 MPa with the OPLS-AA forcefield until they reached the appropriate volume [32]using periodic boundary conditions, a timestep of 1 femtosecond, and a Nose-Hoover thermostat and barostat. Initial velocities were generated using a Maxwell-Boltzmann sampling at 300 K without fixing a random seed. With n = 8, some initialization artifacts still remained, so we computed structure optimization in ASE using a BFGS algorithm and fmax=0.05 with a DFTB+ calculator using a Conjugated Gradient driver to anneal the system.
After the appropriate box volumes were reached, NVT simulations were then conducted at 300K. These production simulations were conducted using DFTB+, initially using a grid of (1, 1, 1) k-points with a Berendsen thermostat, an initial stochastic Maxwell-Boltzmann velocity distribution, a timestep of 0.2 femtoseconds, as well as periodic boundary conditions, and were run until potential energies had converged. For our purposes, an equilibrium cutoff was defined as where the running average of the total energy did not fluctuate by more than 0.1 eV for 2,000 consecutive frames in the NVT simulation. We then selected 1000 frames from the subsequent simulation at random to form testing sets, and used farthest-point sampling[21, 28, 8] on Smooth Overlap of Atomic Positions (SOAP[2]) vectors to select a mutually-exclusive 10,000 frames to constitute our training sets.
Additional testing sets
Decane, dodecane, cyclohexane, 4-propylheptane, and 3,3-diethylpentane, configurations were sampled using the OPLS all-atom (AA) forcefield[32]. Decane and dodecane were simulated with 48 molecules, and the others with 64 molecules, in order to keep computational overhead consistent. The molecules were initialized on a lattice with 10 Å spacing and with each dimension of each lattice position randomly perturbed by -5 to 5 Å sampled from a uniform distribution. We equilibrated the systems with respect to the average radius of gyration, which quantifies the average size of the alkane conformations. The radius of gyration of an alkane, $R_{g}$, is given by $R_{g}=\sqrt{\frac{\sum_{i}m_{i}r_{i}^{2}}{\sum_{i}m_{i}}},$ where the sums are over atoms $i$ in the alkane, $m_{i}$ is the atomic mass, and $r_{i}$ is the displacement relative to the alkane center of mass. We used pyMBAR [12] to calculate the correlation time of the radius of gyration time series. We equilibrated each system for 100 correlation times. Figure S6 demonstrates the equilibrium radius of gyration of the alkanes. Once equilibrated, the systems were run for 1 ns. All classical simulations were run in LAMMPS [47] using forcefield parameters generated by ligpargen[33, 14, 13] in an NPT ensemble at 300 K and 5 MPa using a Nose-Hoover thermostat and barostat with loose couplings of 1 $\text{ps}^{-1}$ each and a timestep of 1 fs.
Energy and force evaluations
For all frames in the training and testing data, we recomputed the energies and forces with DFTB+ using self-consistent charges and the mio Slater-Koster files[17] using a Monkhurst-Pack grid-point interpolation (3, 3, 3). We utilized a high-throughput workflow with signac[1] and signac-flow[46].
To construct values for inter- and intra-molecular energies and forces, we isolated the individual molecules into separate ASE-atoms objects and recomputed energies and forces without consideration of periodic boundary conditions. This yielded
$E_{intermolecular,\ bulk}=E_{total,\ bulk}-E_{intramolecular}$
where $E_{intramolecular}\equiv\sum_{i=1}^{64}E_{i}$, with $E_{i}$ the total energy of the $i^{th}$ single molecule.
Machine-learned potentials and analyses
MACE MLIPs
For every MLIP, we solely used the aforementioned training sets during construction, and trained each model as one normally would for a non-extrapolative task. When training solely on intramolecular quantities (Fig. S3), we down-selected from the 640,000 single-molecule frames to 10,000, choosing one random molecule from each training frame. Testing sets were kept consistent, with 64,000 single molecule frames in total.
The MACE[4] model (Higher‑Order Equivariant Message Passing Neural Network) combines equivariant graph neural network design with explicit many‑body message passing. Unlike most standard MPNNs that rely on only two‑body (pairwise) interactions and require many message‑passing layers to build complexity, MACE constructs higher‑body (up to four‑body) equivariant messages in a hierarchical fashion. This allows the network to achieve high representational power with just two message‑passing iterations, substantially reducing depth, improving parallelism, and maintaining scalability while reaching or surpassing state‑of‑the‑art accuracy on benchmarks such as rMD17, 3BPA, and AcAc[4]. By building hierarchical tensor products of lower‑order features and symmetrizing via generalized Clebsch–Gordan coupling, MACE efficiently implements these many‑body messages without exponential cost. This induces equivariant feature tensors, which are linearly combined (with weights dependent on element type and irreducible representation indices to produce message updates. For our MACE-based potentials, we use the software v.0.3.13 and a standard set of model hyperparameters, as shown in the example script in the Appendix.
(a) Results of predicting per-configuration energies (normalized by the number of atoms) using MACE MLIPs for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
(b) Results of predicting atomic force-vectors using the MACE MLIPs for different train (x-axis) and test (y-axis) pairings, respectively. Errors correspond to the average magnitude of vectorial error across the testing set. Color indicates the mean-absolute-error (MAE) in units of meV/Å, respectively.
(c) The relationship between the shift in composition $(x_{C}^{\text{train}}-x_{C}^{\text{test}})$ and the shift in energies for extrapolated MLIPs. The models trained on the methane, ethane, and propane datasets show a divergence from this proportionality, as they are not necessarily considered “well-conditioned.”
(d) Shifted energy errors for MACE energies, accounting for the learnable shift in Fig. 1(c).
Figure 1: Results of MACE potential building on the total potential energy on n=18n=1-8 alkanes.
SOAP Analysis
Smooth Overlap of Atomic Positions (SOAP) vectors were computed using the featomic[7] and librascal [42] packages with an interaction cutoff of 1010Å (for Fig. 2, Fig. S4) and 77Å (for Fig. 3, Fig. S5). In all cases, we used the GTO basis set with 8 radial components and 4 angular components. Principal Covariates Classification (PCovC, [30]) was computed using v0.3.2 of scikit-matter[21], using as $X$ the aforementioned SOAP vectors of the training set, a mixing parameter of $0.5$, a logistic regression classifier, and oligomer length $n$ as our class value. Each PCovC model was computed solely for like-environments, such as hydrogen, CH2, or tertiary carbon environments (Fig. 2, Fig. S4). Prior to fitting and analysis, feature vectors and energies were appropriately scaled and centered using scikit-matter[21] StandardFlexibleScaler. For regression (Fig. 3), we used scikit-learn[44] RidgeCV with cv=5 and a negative mean squared error scorer. 20 regularization constants ($\alpha$) were considered for RidgeCV using a logarithmic grid.
Results
Extrapolating energies and forces
Fig. 1 shows the results of using trained MLIPs to predict the energies and forces on different oligomeric chain lengths. On first glance, the energy-fitting results of Fig. 1(a) tell a discouraging story – each training set is unable to predict the energies of the other polyalkanes, with a MAE well-above chemical accuracy (which we define as 1meV/atom[53]). However, the force results (Fig. 1(b)) imply something much different, in line with our chemical intuition – at larger chain lengths, we can reasonably extrapolate the energetics of one system to the next (and even those below).
So, why do these two results diverge? A first intuition would be data centering – machine-learning potentials are typically trained to predict the fluctuations in molecular energetics, not their absolute value. In other words, the majority of MLIP methods and software first subtracts the mean of the training targets before determining weights and biases. This means that the predicted energy $E_{pred}$ is ostensibly
$E_{pred}=E_{MLIP}+b,$
where $b$ is a “baselining constant,” or the energetic values shared by all training points. This is verified by the strong colinearity between our predicted and true values, as shown in Fig. S2. A simple explanation of this “mean-shift” would be the shift in atomic energies,
$E_{atom}=n_{C}E_{C}+n_{H}E_{H}=N(x_{C}E_{C}+x_{H}E_{H}),$
which, when learning per-atom energies $E/N$, would yield
$b\equiv(x_{C}^{\text{train}}-x_{C}^{\text{test}})(E_{C}-E_{H}).$
However, this is not the case here, where a shift due solely to atomic contributions would lead to a much larger value than the observed shift. This presents an issue for extrapolating energetics from one system to another, where the only information known a priori may be the difference in underlying chemistry, and where “transfer learning” may be infeasible due to the lack of underlying training data. Thus, we first determine if $b$ is a linear function of our composition ($x_{C}$, $x_{H}$), as this would still allow for the extrapolation of one system to another without prior computation of energetics or configurations.
For models trained on n=48n=4-8 alkanes, there exists a strong linear proportionality (Pearson correlation coefficient = 0.9997, Fig. 1(c)) between the compositional shift and the mean-shift of the datasets. This suggests that these mean-shifts are learnable parameters, even if not obviously grounded in physical values such as $E_{atom}$. To test this, we determined a slope to the line in Fig. 1(c) via non-regularized linear regression, omitting from training the values obtained from models built on methane (blue), ethane (light blue), and propane (green), which are considered unsuccessful models. Accounting for the large co-linearity of the mean-shift and compositional change, many of the error predictions are improved, yet there exist some residual errors (Fig. 1(d)). This suggests that the change in conformational distribution contributes non-negligibly to the mismatch in energies.
However, the results of Fig. 1(b) still demonstrate that there exist two critical lengths of polyalkane for extrapolation: 1) butane, where we see a general reduction in force error magnitude from 20-30meV/Å to 3-6meV/Å (omitting methane), and 2) hexane, after which we see little improvement in force error with additional molecular units. The change in errors from propane to butane follows our chemical intuition – prior to this chain length, there is no sampling of the dihedral rotation possible in these molecules, and thus any system (n=48n=4-8) that contains these interactions would be ill-represented by such models.
But what makes hexane special? To understand this, we can look at the higher-dimensional representation of the interaction environments shared and distinct within the training sets. While we could arguably use the learned features of the MACE MLIPs, we instead employ a preceding, related technology, the Smooth Overlap of Atomic Positions (SOAP[2]), which are deterministic in their description of atomic environments without the implications of training on one dataset versus another . Both MACE and SOAP vectors compute the many-body expansion of atomic interactions, and thus we can infer some behaviors of MACE based on a SOAP representation. We construct our SOAP descriptors based on the distance of the MACE perceptive field (10.0Å).
After we compute the SOAP vectors for each training set, a standard analytical task is to look at the Principal Components Analysis (PCA) of all vectors, which would demonstrate the diversity of environments. However, here we are particularly interested in how the environments within these datasets differ, which is not the underlying mathematical motivation of PCA, and this aspect can be largely lost in PCA (and other unsupervised) maps. So, we instead employ hybrid supervised-unsupervised analyses called Principal Covariates Classification (PCovC)[30, 26, 22, 21], which computes the two-dimensional projection that both retains PCA-like variance and best separates the data based on assigned labels (here the number of carbons in the dataset backbone). Here we focus on the three-body interactions of the CH2 terms, which demonstrate the greatest statistical difference from dataset to dataset, with PCA and PCovC maps of remaining environments in Fig. S4.
Figure 2: Distribution of CH2 environments for the different training sets, mapped via Principal Covariates Classification (PCovC[30]). The axes are directions in latent space that demonstrate the highest degree of diversity in the CH2 environments and separability between the datasets. Points that are overlapping are considered “indistinguishable,” in that an ML model could not tell whether the environment sits in one alkane length or another. The map demonstrates the convergence of environments with increasing $n$. Different alkanes are denoted in different colors (green - propane, yellow - butane, orange - pentane, red - hexane, purple - heptane, and brown - octane), with the corresponding CH2 environment enlarged for each cluster in the map.
From this analysis, we see that the convergence in MLIP error corresponds with the sampling of the CH2 environments three carbons from the end, and the addition of further internal environments within heptane and octane leads to diminishing returns in terms of force approximation. This demonstrates that determining the minimal chain length required for MLIP construction (or molecular fragmentation[20, 29, 41, 48]) amounts to determining the convergence of the majority of molecular environments, a task that can be accomplished via classification models or similar unsupervised tasks on single-molecule data.
Separating the different contributions to total energies
However, predicting the total energies and forces may not be sufficient for a usable potential model, as the hierarchical nature of interactions within these systems decides much of their fundamental behavior[15]. Conceptually, and in classical force fields,, we consider potential energies emerging from three distinct classes of interactions:
$E=E_{atom}+E_{intramolecular}+E_{intermolecular},$
as these terms have significantly different magnitudes, difficulties in calculating, and difficulties in terms of “learnability.” $E_{atom}$ is the largest component but arguably trivial to compute. We can group $E_{atom}$ with intramolecular interactions to obtain an $\tilde{E}_{intramolecular}=E_{atom}+E_{intramolecular}$ by sampling single molecule configurations; this is likewise a large component of $E$. As shown in Fig. S3, extrapolating $\tilde{E}_{intramolecular}$ follows the same lessons learned in predicting $E$ – both in terms of the necessary molecule lengths for extrapolation and the ML baselining leading to divergent narratives for energy- and force-fitting.
The element of this energetic expansion that requires the greatest nuance is the intermolecular interactions $E_{intermolecular}$, which are comparatively weak, longer-range, and computationally expensive to calculate using methods from quantum chemistry[43]. And, when training on total energies, the resultant errors often overshadow these small contributions[15], losing interactions that are decisive in many critical scientific problems.
MACE itself (v.0.3.13) is not natively suited for directly learning these interactions, nor are atom-centered expansions that implicitly or explicitly focus on the nearby atoms in constructing their MLIP[9]. To demonstrate this, we again turn to the SOAP[2] framework, and show how careful attention to weighting interactions can yield better fidelity for intermolecular models. Following the same procedures in terms of training/testing splits, we construct cross-validated Ridge Regressions on SOAP representations for the same n=18n=1-8 $n$-polyalkane datasets used for MACE MLIP construction. SOAP vectors were calculated with a cutoff of 7Å for atomic environments and averaged for each frame such that for the 10,000 training frames and 1,000 testing frames, we have 10,000 and 1,000 corresponding SOAP vectors; we refer to these initial SOAP calculations as the “total” SOAP vectors, or $X^{\text{total}}$.
(a) Regressions using total SOAP vectors $X^{\text{total}}$.
(b) Regressions using the far-sighted SOAP vectors $X^{fs}$.
Figure 3: Results of SOAP-Ridge MLIPs trained on intermolecular potential energy using different SOAP representations.
Similar to the initial results from MACE in Fig. 1(a), the results of using a $E_{\text{intermolecular}}\approx X^{\text{total}}w$, shown in Fig. 3(a), are discouraging. Extrapolation, with a field-standard threshold MAE of 1 meV/atom [53], is only achieved three times (by hexane on heptane, and heptane on hexane/octane), even despite the considerably smaller magnitude of the intermolecular energetics, as compared to the total energies regressed earlier. However, we can adopt lessons from learning the lattice energy of molecular crystals[9] – SOAP representations, and representations that are implicitly or explicitly “near-sighted”[4, 16, 25] will, by design, consider the contributions of nearby atoms more strongly than those further away. This will overshadow the intermolecular environments which contribute to $E_{intermolecular}$, limiting our accuracy in extrapolation.
(a) Force errors resulting from testing n=1-8 trained MLIPs on longer chain lengths and architectures. Color of the bar plot denotes the training set used, and each set of bars refers to a different test molecule.
(b) Distribution of force errors for cyclohexane (left), 4-propylheptane(center), and 3,3-diethylpentane (right), as predicted by a model trained on octane. Colors indicate the atoms within the molecule. For every carbon, the corresponding hydrogens are in the same hue, with a lighter tint.
(c) Intermolecular energy errors resulting from testing n=1-8 trained SOAP-Ridge MLIPs on longer chain lengths and architectures. Color of the bar plot denotes the training set used, and each set of bars refers to a different test molecule. Hatched bars indicate MLIPs constructed using far-sighted SOAP vectors $X^{fs}$, solid bars indicate MLIPs using total SOAP vectors.
Figure 4: Force and energy errors for decane, dodecane, and the nonlinear alkanes.
We can thus construct a “far-sighted” SOAP vector
$X^{fs}=X^{total}-\frac{1}{64}\sum_{i=1}^{64}X_{i}$
where $X_{i}$ is the averaged SOAP vector of the atomic environments in the $i$th molecule within the bulk frame. Effectively, this representation re-weights the feature space by reducing (or eliminating) the influence of contributions from intramolecular environments on the SOAP vector. As shown in Fig. 3(b), using this far-sighted representation results in a significant improvement in the number of extrapolative regimes achieved by SOAP-Ridge MLIPs in comparison to the results in Fig. 3(a). Similar trends are observed as in Fig. 1(b): steep improvements in MAEs are observed between ethane and propane/butane, and once again from pentane to hexane. This follows the convergence in chemical environments (Fig. 2, Fig. S4). Furthermore, these results demonstrate that careful consideration of MLIP features can improve performance for targets other than the total potential energy of a system.
Beyond n=1-8 linear alkanes
It is worth inspecting the limitations of these models when considering longer molecules or more complex hydrocarbon architectures. As shown in Fig. 4(a), decane (n=10) and dodecane (n=12), the trends follow those similar to the shorter oligomers within the series: we see a reduction in error when moving from propane to butane (30meV/Å to 5-6meV/Å), and a saturation near 1.5-2 meV/Å  once training on hexane through octane. This supports our earlier observations considering the saturation of interaction environments, and suggests that these trends should continue to longer linear alkanes. Returning to the SOAP-Ridge MLIPs constructed using total SOAP vectors ($X^{total}$) and far-sighted SOAP vectors ($X^{fs}$) in Fig. 4(c), we also maintain previous trends: intermolecular energy MAEs decrease consistently as the training alkane chain length increases, and MLIPs constructed using $X^{fs}$ consistently outperform MLIPs constructed with $X^{total}$. This re-confirms our earlier claim that re-weighting the feature-space to exclude intramolecular contributions enables higher accuracy when targeting intermolecular energy.
Molecules of different architectures fare less well in extrapolation. For cyclohexane, 4-propylheptane, and 3,3-diethylpentane, we typically converge to higher values (6-10meV/Å) than within the linear series (1-2meV/Å). On one hand, this is due to the lack of longer alkyl chains within cyclohexane, 4-propylheptane, and 3,3-diethylpentane – moving to hexane, heptane, and octane training data is not nearly as beneficial when the longest uninterrupted alkyl chain is three carbons long. For 4-propylheptane and 3,3-diethylpentane, the larger errors are largely attributed to the tertiary and quaternary carbons (Fig. 4(b)), neither of which exists in a linear alkane. Though we see higher errors from the MACE predicted forces for these molecules, we maintain consistent trends in intermolecular energies as predicted by SOAP-Ridge MLIPs with decreasing errors as training alkane chain length increases, and lower errors for all but one far-sighted MLIPs.
Cyclohexane presents a more interesting case. The carbons in cyclohexane have long received different parameterization than those in hexane via popular force fields such as OPLS-AA[34, 18, 32]. This has largely been attributed to the constrained torsion angles in cyclohexane, which remove the freely rotating C-C dihedrals that characterize linear alkanes[34]. However, here, the key difference is the local environment distribution: linear alkanes provide very few CH2 environments with neighboring CH2 at the 232-3Å range, whereas cyclohexane contains several such neighbors because of its compact ring geometry. As a result, when the model trained only on octane is evaluated on cyclohexane, it encounters CH2–CH2 configurations that were poorly sampled in the training set, producing the shift observed in Fig. 4(b).
Conclusions
In this study, we have demonstrated key behaviors of MLIPs that can be applied to build predictive models for oligomeric systems, with implications on MLIP extrapolation for macromolecular and polymeric materials. First, we demonstrate and visualize that, in systems with similar chemistries, extrapolation of forces and energies is possible, provided that those environments present in the target molecule are properly sampled within the smaller analogs. Specifically for alkanes, we see a strong reduction in error with butane, and we see the force errors converging once we reach hexane. These results were also demonstrated on longer linear alkanes and different architectures. The latter proved more nuanced and complicated as an extrapolation task, as while non-linear alkanes contain the same chemistry, MLIPs do not consider these aspects explicitly, but rather as emerging from data distributions. And, because the environmental distributions of branched or cyclical alkanes are fundamentally different from linear alkanes, the MLIPs fare worse in extrapolating to these systems.
Second, we showed that extrapolation of energies results in a learnable shift between different molecules; this can be largely ameliorated by predicting the mean shift in probability distributions as a function of composition. As forces are not affected by constant shifts in energy, the extrapolation of forces between different molecules is straightforward and does not require accounting for this effect.
Finally, we demonstrated a powerful lesson about the hierarchy of energies. Intermolecular contributions are important in determining thermodynamic properties, but have historically been challenging to predict due to their small magnitude relative to the intramolecular component. We have shown that careful partitioning of molecular descriptors can yield MLIPs capable of predicting and extrapolating intermolecular energetics. Given the relative difficulty of computing intra- versus intermolecular energetics, this suggests that the most difficult calculations, those relating the interactions of one molecule to another, are the easiest to translate from one system to another of similar chemistry.
Research Funding
This project was funded by the Wisconsin Alumni Research Fund (RKC) and by NSF through the University of Wisconsin Materials Research Science and Engineering Center (DMR-2309000, AL, CC). NH acknowledges support from the Michael and Virginia Conway Professorship. The funders played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
Data Availability
The datasets generated and/or analysed during the current study are available in the Zenodo repository at [27] [DOI: 10.5281/zenodo.17227874].
Code Availability
The underlying code used in this study are available via open-source software through GitHub, namely MACE (https://github.com/ACEsuit/mace[4]), scikit-matter (https://github.com/scikit-learn-contrib/scikit-matter[22]), featomic (https://github.com/metatensor/featomic[7]), and librascal (https://github.com/lab-cosmo/librascal.git[42]).
Author Contributions
Project conceptualization (NH, RC), data curation (NH, CC, AL), formal analysis (NH, RC), methodology (NH, RC), supervision (AL, RC), writing (NH, AL, CC, RC).
Competing Interests
All authors declare no financial or non-financial competing interests.